OpenStudy (anonymous):

I am trying to solve this problem. A car traveling with an initial velocity of 13 m/s accelerates at a constant rate of 2.0 m/s2 for a time of 2 seconds.

OpenStudy (anonymous):

are you trying to find the final velocity?

OpenStudy (anonymous):

i found the final velocity, now i am working on calculating the distance during that time.

OpenStudy (potatoes.ramu):

Use the equation s=ut+1/2(at^2)

OpenStudy (potatoes.ramu):

Where s is the distance u is the initial velocity v is the final velocity a is the acceleration and t is the time :)

OpenStudy (potatoes.ramu):

Or you can use the equation v^2 = u^2 + 2as

OpenStudy (potatoes.ramu):

To find the distance :)